OpenStudy (anonymous):

HELP!? Solve the absolute value equation. Graph the solution. 2|3x – 5| – 8 = 8

OpenStudy (anonymous):

2I3x-5I-8=8 2I3x-5I=8+8 2I3x-5I=16 I3x-5I=8

OpenStudy (anonymous):

understood it?

OpenStudy (anonymous):

up to there?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

the answers have an answer than or and than another answer like: 51/4 or -51/4

OpenStudy (anonymous):



OpenStudy (anonymous):

i) x>3/5 => 3x-5=8 x=13/3 ii)x<3/5 => -(3x-5)=8 -3x = 3 x=-1 X1 = 13/3 X2= -1